# CVE-Bench: A Benchmark for AI Agents’ Ability to Exploit Real-World Web Application Vulnerabilities

### v1

@@ Abstract
Large language model (LLM) agents are increasingly capable of autonomously conducting cyberattacks, posing significant threats to existing applications. This growing risk highlights the urgent need for a real-world benchmark to evaluate the ability of LLM agents to exploit web application vulnerabilities. However, existing benchmarks fall short as they are limited to abstracted Capture the Flag competitions or lack comprehensive coverage. Building a benchmark for real-world vulnerabilities involves both specialized expertise to reproduce exploits and a systematic approach to evaluating unpredictable threats. To address this challenge, we introduce CVE-Bench, a real-world cybersecurity benchmark based on critical-severity Common Vulnerabilities and Exposures. In CVE-Bench, we design a sandbox framework that enables LLM agents to exploit vulnerable web applications in scenarios that mimic real-world conditions, while also providing effective evaluation of their exploits. Our evaluation shows that the state-of-the-art agent framework can resolve up to 13% of vulnerabilities.
@@ 1 Introduction
In recent years, large language model (LLM) agents have increasingly demonstrated capabilities in complex tasks that require reasoning (Jaech et al., 2024) and tool use (Wu et al., 2024), including resolving GitHub issues (Yang et al., 2024a; Jimenez et al., 2023), fixing bugs (Mündler et al., 2024), and interacting with real computing environment (Xie et al., 2024). The advancement of these capabilities has raised concerns about the potential misuse of LLM agents in conducting cyberattacks (Abdali et al., 2024). Consequently, there has been an increasing effort from government agencies (Raimondo, 2024), industry practitioners (Hurst et al., 2024), and researchers (Fang et al., 2024a, c; Abdali et al., 2024; Zhou et al., 2024; Yang et al., 2024b; Guo et al., 2024; Zhang et al., 2024) to evaluate and red-team with LLM agents. This effort is particularly critical for web applications, which are prime targets for cyberattacks due to their importance as entry points to vital services and repositories of sensitive user data (Huang et al., 2017; OWASP, 2021). For example, a vulnerability in Twitter’s system resulted in significant data breaches affecting over 5.5 million people from 2014 to 2020 (Twitter, 2022; Winde, 2022).
Unfortunately, existing benchmarks do not adequately evaluate the capability of LLM agents to exploit real-world vulnerabilities of web applications. These benchmarks focus on short code-snippets (Zhou et al., 2024) or abstracted “Capture The Flag” (CTF) challenges (Zhang et al., 2024; Yang et al., 2023). In contrast, exploiting real-world vulnerabilities introduces more complexity that requires not only interacting with the web application, but also understanding the application architecture and executing attacks that could affect the web server or its users. Furthermore, previous research assessing the abilities of LLM agents to exploit real-world vulnerabilities offers only a limited range of tasks and attack types, which are insufficient to simulate a production scenario effectively (Fang et al., 2024a, c).
@@
Overcoming the limitation of prior work and building a real-world cybersecurity benchmark is especially challenging. First, ensuring comprehensive coverage requires setting up a wide variety of vulnerable web applications and guaranteeing that their vulnerabilities are reproducible. Second, to ensure the correctness of the benchmark, we must provide reference exploits. Manually exploiting a vulnerability can be complicated and requires an in-depth understanding of web architecture, analyzing the vulnerability and corresponding patches (if any), identifying security weaknesses, and devising feasible exploits to compromise the application. Such a process is notoriously time-consuming and labor-intensive (Mu et al., 2018), costing 5-24 person-hours to reproduce and exploit a single vulnerability in our benchmark. Finally, to rigorously access whether an exploit by LLM agents has been successful, we need to effectively detect attacks for each web application. Unfortunately, cyberattack detection is a long-standing research problem, requiring sophisticated strategies and lacking a one-size-fit-all solution (Raiyn et al., 2014; Singh & Silakari, 2009; Ahmetoglu & Das, 2022).
We address these challenges through a systematic sandbox framework that makes a real-world cybersecurity benchmark feasible (Figure 1). For each vulnerability, we implement a collection of containers (i.e., target containers) designed to host web application with exposed vulnerabilities. To evaluate the diverse strategies LLM agents might use to exploit vulnerabilities, we standardize potential attack vectors into eight standard attacks and develop an evaluation system to automatically grade LLM agents. Then, the agents are directed to achieve any one of the eight standard attack targets. In addition, to ensure vulnerabilities in the benchmark are exploitable, we implemented or reproduced a reference exploit for each vulnerability as a proof of concept.
@@
We apply CVE-Bench to evaluate various LLM agents under both zero-day and one-day settings. Our findings indicate that existing LLM agents designed for cybersecurity, such as the agents developed in Cybench (Zhang et al., 2024), exhibit significant shortcomings, achieving a success rate of 2.5% with five attempts in the one-day setting. Furthermore, when equipped with appropriate tools, such as sqlmap (sqlmap, 2024), teams of LLM agents demonstrate substantial improvement, achieving a success rate as high as 13% with five attempts in the one-day setting.
@@ 2 Background
Existing LLM Agents for Cyberattacks. Prior work has sought to analyze the cybersecurity threats introduced by the development of LLMs via designing various agent frameworks for conducting cyberattacks. Instead of directly prompting LLMs, Cybench proposed an agent framework that uses loops of actions: act, execute, and update, to effectively analyze feedback from the environment (Zhang et al., 2024). This reactive approach, or ReAct-style agent frameworks (Yao et al., 2023), has also applied to exploit vulnerabilities in web applications based on known vulnerability descriptions, commonly referred to as the one-day setting (Fang et al., 2024b, a). More recently, agent teams with hierarchical planning and task-specific agents have been developed to hack web applications under the zero-day setting (Fang et al., 2024c). This framework consists of teams of specialized hacker agents, each an expert in a specific cybersecurity area such as cross-site scripting (XSS) or SQL injection, and supervisor agents responsible for strategic planning and directing the hacker agents. These agents frameworks highlight the significant threats posed by using LLMs to autonomously execute cyberattacks. Thus, real-world cybersecurity benchmarks are crucial for comprehensive evaluation.
Existing Cybersecurity Benchmarks are Insufficient.
Features Cybench Fang et al. (2024a; 2024c) CVE-Bench
@@
[…]
As shown in Table 1, existing benchmarks for evaluating LLM agents in cybersecurity have several limitations. Although Cybench includes a significant amount of vulnerabilities, its vulnerabilities and tasks do not reflect real-world scenarios, focusing instead on vulnerabilities of smaller databases without severity ratings (Zhang et al., 2024). Furthermore, Cybench is limited to CTF tasks and neglects to evaluate other severe attacks, such as database modification. More recently, Fang et al. (2024a, c) built benchmarks that involve medium-to-critical real-world CVEs across a variety of attack types. However, they only include a limited number of vulnerabilities and evaluate just one specific attack type per CVE. In contrast, our benchmark matches Cybench in scale while incorporating real-world, critical-severity vulnerabilities and supporting a diverse range of attack types.
We introduce the details of our benchmark, CVE-Bench, [^1] Data and code are available at https://github.com/uiuc-kang-lab/cve-bench.git. in the following section.
@@ 3 CVE-Bench
We present an overview of CVE-Bench, followed by details of task specification and the benchmark construction process. We describe one of the tasks as an example.
@@ 3.1 Overview
We built CVE-Bench based on web-related CVEs from the National Vulnerability Database (NVD), a repository of software vulnerability maintained by the U.S. government (Booth et al., 2013). Matching the scale of Cybench (Zhang et al., 2024), CVE-Bench includes 40 CVEs. As shown in Table 2, CVEs in CVE-Bench have a minimum of 9.0 severity score measured by the Common Vulnerability Scoring System (CVSS) version 3.1 base score. We further categorize CVEs based on type of the vulnerable web applications in Table 3. As shown, we include a wide range of application types, with a focus on content management, such as WordPress (WordPress, 2011), and AI applications, such as LoLLMs Aloui (2025).
@@
Table 2: Distribution of based severity scores of CVEs in CVE-Bench. The severity score is calculated according to the base score of the Common Vulnerability Scoring System (CVSS) version 3.1.
@@ 3.2 Task Specification
A Denial-of-Service (DoS) attack is a critical cyber threat that makes a service temporarily or indefinitely unavailable to its intended users. Historically, DoS attacks have resulted in substantial financial losses, amounting to millions of dollars (Loukas & Öke, 2010). In CVE-Bench, we use DoS as the first standard attack.
@@
Vulnerabilities of web applications can potentially allow unauthorized access to the file system of a host machine, resulting in severe consequences such as data leaks and file system overloading. Therefore, we select file access and file creation as two of our standard attacks. To demonstrate successful file access, LLM agents need to retrieve and send both the name and content of a file to a designated endpoint, where our grader will verify the accuracy of the file content. For file creation, the LLM agents must create the file named /tmp/pwned, a location that should be inaccessible to application users.
@@
Using a separate database management system is a convenient and efficient way for web developers to manage data in web applications. Unfortunately, databases can become a significant attack targets, leading to vast data breaches of sensitive user data. Historically, various strategies, such as privilege escalation and SQL injection, are used to compromise the database, leading to significant financial loss and affecting millions of individuals (U.S. Department of Justice, 2009). Thus, we consider database modification and access as two standard attacks. For database modification, the LLM agents need to change the data as much as possible. To exploit a database access vulnerability, the LLM agents need to retrieve the data in a specific table.
@@
Login interfaces of web applications can also be vulnerable, allowing attackers to gain unauthorized administrator access or escalate the privileges of low-privilege users. Such attacks can cause severe consequences, such as complete service takeover and data breach (Thomas et al., 2017). Thus, we consider unauthorized administrator login and privilege escalation as two standard attacks. To achieve unauthorized administrator login, LLM agents need to log into the application using a provided administrator username without authorization. For privilege escalation, agents must enhance the privileges of a specified user account.
@@
Disallowed Attacks. To keep LLM agents focused on target attacks, we specify a list of out-of-scope attacks that are disallowed. We only allow attacks on the given websites and nothing else. For example, attacking the evaluator is prohibited. We also specify that LLM agents not hack passwords by brute force. To discourage such attacks, we use random generated string as passwords for administrative accounts, database users, and database root users.
@@
We defer examples of complete task specifications to Appendix B.
@@ 3.3 Benchmark Construction
We built CVE-Bench based on CVEs from the NVD. For each CVE, we implement a set of containers to host the vulnerable web application, reproduce the exploits to demonstrate the existence of the vulnerability, and grade the LLM agents.
@@
Across the selected CVEs, 24 of them have login interfaces and use separate databases. We defer a complete list of CVEs and their detailed information to Appendix A.
@@ 3.4 Task Example
Reproduced Exploit. We host our exploit reproduction in a separate container that is connected to the target containers. Based on the published patch for this vulnerability, we located the vulnerability at the endpoint: /wp-admin/admin.php?page=tutor. Then, we craft a payload that bypass the authentication of WordPress and delete data entries in the database.
@@
    Inspect the received content in the proof upload endpoint to check if LLM agents uploaded correct content of the table secret.
@@
    Query the last login time of the administrator to check LLM agents successfully logged in as an administrator.
@@
    Check whether the privilege of the default user is escalated to administrators.
@@ 4.1 Settings
    Cybench Agent or Cy-Agent (Zhang et al., 2024): Cy-Agent is developed as the default LLM agent framework for Cybench. In each iteration, it applies an LLM to decide an action based on the memory, execute the action in the environment, and update the memory based on the observation of the environment.
@@
    AutoGPT (Richards, 2023): AutoGPT is a general agent framework designed for automating complex workflow with LLMs. It enables LLM with planning and tool using. In each iteration, AutoGPT first summarizes the observation, and then reasons, self-criticizes, and plans the next step. During execution, AutoGPT chooses the proper tool to use.
We defer detailed configurations and prompts of these Agents to Appendix C.
@@ 4.2 Results
Figure 3: Success rates of different LLM agents on CVE-Bench. LLM agents can expoit up to 13% and 25% vulnerabilities under zero-day and one-day settings, respectively.
@@
Table 4: Per-task costs of evaluating LLM agents on CVE-Bench.
Success Rate. We present success rates of different LLM agents in Figure 3 with one or five attempts. As shown, LLM agents can exploit up to 13% web application vulnerabilities under the zero-day setting, and 25% under the one-day setting. Except for AutoGPT, agents generally achieved higher success rates under one-day setting than the zero-day setting, since more relevant information (i.e., vulnerability descriptions) is provided under the one-day setting.
We observe that AutoGPT demonstrates superior performance by achieving the highest success@5 rate, with an unexpectedly higher zero-day success@5 rate compared to its one-day success@5 rate. Upon reviewing the reasoning logs, we find that under the zero-day setting, AutoGPT could identify and exploit vulnerabilities that are easier than those provided in the one-day description. We analyzed one of such cases in Section 4.3 (CVE-2024-37831).
@@
Figure 4: Distribution of successful exploits by Cy-Agent, T-Agent, and AutoGPT.
Exploit Composition. To understand why T-Agent and AutoGPT achieves good performance, we take a deeper inspection and present the composition of successful exploits in Figure 4. As shown, among successful exploits, T-Agent performs 68% and 30% database access under zero-day and one-day settings, respectively, while the percentage of database access is smaller for AutoGPT: 0% in the both zero-day and one-day settings. This is because T-Agent is equipped with sqlmap to help with perform SQL injection attacks. With sqlmap, T-Agent can locate the vulnerability and perform SQL injection automatically. According to the reasoning traces, Cy-Agent attempts sqlmap for most of the CVEs. Appropriate use of sqlmap can significant improve the success rate of exploit SQL injection vulnerabilities, the second most frequent vulnerability (OWASP, 2021), which can lead to unauthorized database access and data breaches.

### v4

@@ Abstract
Large language model (LLM) agents are increasingly capable of autonomously conducting cyberattacks, posing significant threats to existing applications. This growing risk highlights the urgent need for a real-world benchmark to evaluate the ability of LLM agents to exploit web application vulnerabilities. However, existing benchmarks fall short as they are limited to abstracted Capture-the-Flag competitions or lack comprehensive coverage. Building a benchmark for real-world vulnerabilities involves both specialized expertise to reproduce exploits and a systematic approach to evaluating unpredictable attacks. To address this challenge, we introduce CVE-Bench, a real-world cybersecurity benchmark based on critical-severity Common Vulnerabilities and Exposures. In CVE-Bench, we design a sandbox framework that enables LLM agents to exploit vulnerable web applications in scenarios that mimic real-world conditions, while also providing effective evaluation of their exploits. Our experiments show that the state-of-the-art agent framework can exploit up to 13% of the vulnerabilities.
@@ 1 Introduction
In recent years, large language model (LLM) agents have increasingly demonstrated capabilities in complex tasks that require reasoning (Jaech et al., 2024) and tool use (Wu et al., 2024), including resolving GitHub issues (Yang et al., 2024a; Jimenez et al., 2023), fixing bugs (Mündler et al., 2024), and interacting with real computing environments (Xie et al., 2024). The advancement of these capabilities has raised concerns about the potential misuse of LLM agents in conducting cyberattacks (Abdali et al., 2024). Consequently, there has been increasing efforts from government agencies (Raimondo, 2024), industry practitioners (Hurst et al., 2024), and researchers (Fang et al., 2024a, c; Abdali et al., 2024; Zhou et al., 2024; Yang et al., 2024b; Guo et al., 2024; Zhang et al., 2024a) to evaluate and red-team with LLM agents. This effort is particularly critical for web applications, which are prime targets for cyberattacks due to their importance as entry points to vital services and repositories of sensitive user data (Huang et al., 2017; OWASP, 2021). For example, a vulnerability in Twitter’s system resulted in significant data breaches affecting over 5.5 million people from 2014 to 2020 (Twitter, 2022; Winde, 2022).
Unfortunately, existing benchmarks do not adequately evaluate the capabilities of LLM agents to exploit real-world vulnerabilities of web applications. These benchmarks focus on short code-snippets (Zhou et al., 2024) or abstracted “Capture The Flag” (CTF) challenges (Zhang et al., 2024a; Yang et al., 2023; Shao et al., 2024; Bhatt et al., 2024; Wan et al., 2024). In contrast, exploiting real-world vulnerabilities introduces more complexity that requires not only interacting with the web application, but also understanding the application architecture and executing attacks that could affect the web server or its users. Furthermore, previous research assessing the abilities of LLM agents to exploit real-world vulnerabilities offers only a limited range of tasks and attack types, which are insufficient to simulate a production scenario effectively (Fang et al., 2024a, c).
@@
Overcoming the limitation of prior work and building a real-world cybersecurity benchmark is especially challenging. First, ensuring comprehensive coverage requires setting up a wide variety of vulnerable web applications and guaranteeing that their vulnerabilities are reproducible. Second, to ensure the correctness of the benchmark, we must provide reference exploits. Manually exploiting a vulnerability can be complicated and requires an in-depth understanding of web architecture, analyzing the vulnerability and corresponding patches (if any), identifying security weaknesses, and devising feasible exploits to compromise the application. Such a process is notoriously time-consuming and labor-intensive (Mu et al., 2018), costing 5-24 person-hours to reproduce and exploit a single vulnerability in our benchmark. Finally, to rigorously assess whether exploits by LLM agents are successful, we should be able to detect any forms of cyberattack for all web applications. Unfortunately, cyberattack detection is a long-standing research problem, requiring sophisticated strategies and lacking a one-size-fits-all solution (Raiyn et al., 2014; Singh & Silakari, 2009; Ahmetoglu & Das, 2022).
We address these challenges through a systematic sandbox framework that makes a real-world cybersecurity benchmark feasible (Figure 1). For each vulnerability, we implement a collection of containers (i.e., target containers) designed to host a web application with exposed vulnerabilities. To evaluate the diverse strategies LLM agents might use to exploit vulnerabilities, we standardize potential attack vectors into eight standard attacks and develop an evaluation system to automatically grade LLM agents. Then, the agents are directed to achieve any one of the eight standard attack targets. In addition, to ensure vulnerabilities in the benchmark are exploitable, we reproduce a reference exploit for each vulnerability as a proof of concept.
@@
We apply CVE-Bench to evaluate various LLM agents under both zero-day and one-day settings. Our findings indicate that existing LLM agents designed for cybersecurity, such as the agent developed in Cybench (Zhang et al., 2024a), exhibit significant shortcomings, achieving a success rate of 2.5% with five attempts in the one-day setting. Furthermore, with a hierarchical multi-agent framework, teams of LLM agents (Fang et al., 2024c) demonstrate substantial improvement, achieving a success rate as high as 13% with five attempts in the one-day setting.
@@ 2 Background
Existing LLM Agents for Cyberattacks. Prior work has sought to analyze the cybersecurity threats introduced by the development of LLMs via designing various agent frameworks for conducting cyberattacks. Instead of directly prompting LLMs, Cybench proposed an agent framework that uses loops of actions: act, execute, and update, to effectively analyze feedback from the environment (Zhang et al., 2024a). This reactive approach, or ReAct-style agent frameworks (Yao et al., 2023), has also been applied to exploit vulnerabilities in web applications based on known vulnerability descriptions, commonly referred to as the one-day setting (Fang et al., 2024b, a). More recently, agent teams with hierarchical planning and task-specific agents have been developed to hack web applications under the zero-day setting (Fang et al., 2024c). This framework consists of teams of specialized hacker agents, each an expert in a specific cybersecurity area such as cross-site scripting (XSS) or SQL injection, and supervisor agents responsible for strategic planning and directing the hacker agents. These agentic frameworks highlight the significant threats posed by using LLMs to autonomously execute cyberattacks. Thus, real-world cybersecurity benchmarks are crucial for comprehensive evaluation.
[…]
Features Cybench (2024a) Fang et al. (2024a; 2024c) CVE-Bench
@@
Existing Cybersecurity Benchmarks are Insufficient. As shown in Table 1, existing benchmarks for evaluating LLM agents in cybersecurity have several limitations. Although existing CTF-based benchmarks include a significant amount of vulnerabilities, their vulnerabilities and tasks do not reflect real-world scenarios, focusing instead on vulnerabilities of smaller databases without severity ratings (Zhang et al., 2024a; Yang et al., 2023; Shao et al., 2024; Bhatt et al., 2024; Wan et al., 2024). Furthermore, these benchmarks are limited to CTF tasks and neglect to evaluate other severe attacks, such as database modification. Recently, Fang et al. (2024a, c) built benchmarks that involve medium-to-critical real-world CVEs with various attack types. However, they only include a limited number of vulnerabilities and evaluate just one specific attack type per CVE. In contrast, our benchmark matches Cybench in scale while incorporating real-world, critical-severity vulnerabilities and supporting a diverse range of attack types. Recent work has proposed benchmarks for the security of AI systems (Zhan et al., 2024; Zhang et al., 2024b), which are orthogonal to our work.
We introduce the details of our benchmark, CVE-Bench, [^1] Data and code are available at https://github.com/uiuc-kang-lab/cve-bench. in the following section.
@@ 3 CVE-Bench
We present an overview of CVE-Bench, followed by details of task specification and the benchmark construction process. We describe one of our tasks as an example.
@@ 3.1 Overview
We build CVE-Bench based on web-related CVEs from the National Vulnerability Database (NVD), a repository of software vulnerability maintained by the U.S. government (Booth et al., 2013). Matching the scale of Cybench (Zhang et al., 2024a), CVE-Bench includes 40 CVEs. As shown in Table 2, CVEs in CVE-Bench have a minimum of 9.0 severity score measured by the Common Vulnerability Scoring System (CVSS) version 3.1 base score. We further categorize CVEs based on type of the vulnerable web applications in Table 3. As shown, we include a wide range of application types, with a focus on content management, such as WordPress (WordPress, 2011), and AI applications, such as LoLLMs Aloui (2025).
Table 2: Distribution of base severity scores of CVEs in CVE-Bench. The severity score is calculated according to the base score of the Common Vulnerability Scoring System (CVSS) version 3.1.
@@
Table 3: Distribution of types of web applications in CVE-Bench.
@@ 3.2 Task Specification
A Denial-of-Service (DoS) attack is a critical cyber threat that makes a service temporarily or indefinitely unavailable to its intended users. Historically, DoS attacks have resulted in substantial financial losses, amounting to millions of dollars (Loukas & Öke, 2010). In CVE-Bench, we use DoS as a standard attack.
@@
Vulnerabilities of web applications can potentially allow unauthorized access to the file system of a host machine, resulting in severe consequences such as data leaks and file system overloading. Therefore, we select file access and file creation as two of our standard attacks. To demonstrate successful file access, LLM agents need to retrieve and send both the name and content of a file to a designated endpoint, where our grader will verify the accuracy of the file content. For file creation, the LLM agents must create a file at /tmp/pwned, a location that should be inaccessible to application users.
@@
Using a separate database management system is a convenient and efficient way for web developers to manage data in web applications. Unfortunately, databases can become a significant attack target, leading to vast breaches of sensitive user data. Historically, various strategies, such as privilege escalation and SQL injection, are used to compromise the database, leading to significant financial loss and affecting millions of individuals (U.S. Department of Justice, 2009). Thus, we consider database modification and access as two standard attacks. For database modification, the LLM agents need to change the data as much as possible. To exploit a database access vulnerability, the LLM agents need to retrieve the data in a specific table.
@@
Login interfaces of web applications can also be vulnerable, allowing attackers to gain unauthorized administrator access or escalate the privileges of low-privilege users. Such attacks can cause severe consequences, such as complete service takeover (Thomas et al., 2017). Thus, we consider unauthorized administrator login and privilege escalation as two standard attacks. To achieve unauthorized administrator login, LLM agents need to log into the application using a provided administrator username without authorization. For privilege escalation, agents must enhance the privileges of a specified user account.
@@
Disallowed Attacks. To keep LLM agents focused on target attacks, we specify a list of out-of-scope attacks that are disallowed. We only allow attacks on the given websites and nothing else. For example, attacking the evaluator is prohibited. We also specify that LLM agents not hack passwords by brute force. To discourage such attacks, we use randomly generated strings as passwords for administrative accounts, database users, and database root users.
@@
We defer examples of complete task specifications to Appendix C.
@@ 3.3 Benchmark Construction
We built CVE-Bench based on CVEs from the NVD. For each CVE, we (i) containerized the vulnerable web application, (ii) reproduced the reference exploit to verify the validity of our implementation, and (iii) developed an evaluation server that automatically determines whether an LLM agent has succeeded.
@@
Across the selected CVEs, 24 of them have login interfaces and use separate databases. We defer a complete list of CVEs and a detailed description of our construction process to Appendix A.
@@ 3.4 Task Example
Reproduced Exploit. We host our exploit reproduction in a separate container that is connected to the target containers. Based on the published patch for this vulnerability, we located the vulnerability at the endpoint: /wp-admin/admin.php?page=tutor. Then, we craft a payload that bypasses the authentication of WordPress and deletes data entries in the database.
@@
    Inspect the received content in the proof-uploading endpoint to check if LLM agents uploaded correct content of the table secret.
@@
    Query the last login time of the administrator to check if LLM agents successfully logged in as an administrator.
@@
    Check whether the privilege of the default user is escalated to an administrator.
@@ 4.1 Settings
    Cybench Agent or Cy-Agent (Zhang et al., 2024a): Cy-Agent is an agent used in Cybench for cybersecurity challenges. In each iteration, it applies an LLM to decide an action based on the memory, execute the action in the environment, and update the memory based on the observation of the environment.
@@
    AutoGPT (Richards, 2023): AutoGPT is a general agent framework designed for automating complex workflow with LLMs. It enables LLM to plan actions and use tools. In each iteration, AutoGPT first summarizes the observation, and then reasons, self-criticizes, and plans the next step. During execution, AutoGPT chooses the proper tool to use.
We framed tasks in an ethical context such that none of the agents refused the exploitation requests. In the prompt, we instruct agents to act as white-hat hackers with permissions granted by application owners. We defer detailed configurations and prompt templates to Appendix C and other baselines to Appendix D.
@@ 4.2 Results
Figure 3: Success rates of different LLM agents on CVE-Bench. LLM agents can exploit up to 10% and 13% vulnerabilities under zero-day and one-day settings, respectively.
Table 4: Per-task costs of evaluating LLM agents on CVE-Bench.
@@
Success Rate. We present success rates of different LLM agents in Figure 3 with one or five attempts. As shown, LLM agents can exploit up to 10% web application vulnerabilities under the zero-day setting, and 12.5% under the one-day setting. Except for AutoGPT, agents generally achieved higher success rates under one-day setting than the zero-day setting, since more relevant information (i.e., vulnerability descriptions) is provided under the one-day setting.
We observe that AutoGPT demonstrates superior performance by achieving the highest success@5 rate, with an unexpectedly higher zero-day success@5 rate compared to its one-day success@5 rate. Upon reviewing the reasoning logs, we find that under the zero-day setting, AutoGPT could identify and exploit new vulnerabilities that are easier than those provided in the one-day description. We analyzed one of such cases in Section 4.3 (CVE-2024-37831).
@@
Figure 4: Distribution of successful exploits by Agents. We only show the types of attack conducted successfully.
Exploit Composition. To understand why T-Agent and AutoGPT achieved good performance, we take a deeper inspection and present the composition of successful exploits in Figure 4. As shown, among successful exploits, T-Agent performs 68% and 30% database access under zero-day and one-day settings, respectively, while the percentage of database access is smaller for AutoGPT: 0% in the both zero-day and one-day settings. This is because T-Agent does better in using sqlmap to perform SQL injection attacks. With a multi-agent framework, T-Agent can strategically plan and execute a complete SQL injection workflow, including the localization of vulnerability endpoints and testing of different inject techniques. Appropriate use of sqlmap can significant improve the success rate of exploit SQL injection vulnerabilities, the second most frequent vulnerability (OWASP, 2021), which can lead to unauthorized database access and data breaches.